# Out-of-Distribution Detection: A Task-Oriented Survey of Recent Advances

### v3

@@ 5.1. Zero-shot OOD Detection Approaches
where $x’$ ’ denotes the test image, $Y_{in}$ is the set of in-distribution labels, $T(t_{i})$ is the text embedding for prompt $t_{i}$ , $s_{i}(x’)$ ’ ) is the cosine similarity between the image and text features, and $\tau$ is the softmax temperature.
@@
Recent works have further enriched the textual context used by CLIP to improve OOD detection. For example, NegLabel (Anonymous, 2024), LAPT (Zhang et al., 2024a), CSP (Chen et al., 2024b), and CLIPScope (Fu et al., 2024) all leverage curated corpora to augment the set of candidate labels. Specifically, NegLabel selects negative labels from a corpus that are semantically dissimilar to ID classes, thereby sharpening the model’s ability to distinguish OOD samples. CLIPScope (Fu et al., 2024) goes a step further by integrating these negative labels into a Bayesian scoring framework to revise the original OOD score. LAPT, in contrast, utilizes text-to-image generation or image retrieval models to obtain visual exemplars for both ID and negative labels, followed by prompt tuning to optimize OOD discrimination. It is worth emphasizing that methods such as ZOC (Esmaeilpour et al., 2022), CLIPN (Wang et al., 2023c), NegLabel (Anonymous, 2024), and CLIPScope (Fu et al., 2024) each propose distinct OOD scoring mechanisms tailored to their respective architectures. In contrast, the MCM score stands out for its generalizability and can be seamlessly integrated into diverse CLIP-based frameworks.
Recent advances in zero-shot OOD detection have moved beyond static negative sampling by leveraging more adaptive strategies within the joint vision-language space. For example, AdaNeg (Zhang and Zhang, 2024) dynamically generates negative proxies at test time to better match the semantics of OOD samples, while AdaND (Cao et al., 2025) focuses on mitigating the impact of noisy samples by decoupling noise detection from classification using CLIP-based pseudo-labels. In contrast, SimLabel (Zou et al., 2025) distinguishes OOD inputs by analyzing the consistency of image-label similarities among semantically related classes, automatically constructing similar class groups via hierarchical structures and large language models. These approaches thus differ in their strategies for negative sampling, noise handling, and semantic grouping, reflecting the diverse directions in recent OOD detection research.
Complementing these proxy- and consistency-based advances, several works have proposed to enhance OOD detection by expanding the diversity and expressiveness of negative or outlier labels. OLE (Ding and Pang, 2024) exposes CLIP to a large set of outlier labels, selected from external corpora and clustered into prototypes, to better model the outlier space. Furthermore, OLE augments these prototypes through interpolation with ID representations, effectively filling the semantic gap between ID and OOD regions in the embedding space. Similarly, recent work on prompt learning for near-OOD detection (Jung et al., 2024) introduces novel scoring functions, such as the Empty-Class Score, which measures the alignment between an image and the context vector excluding any class name, and relative scores that subtract this alignment from standard logit-based OOD scores. This simple post-hoc adjustment provides substantial gains, especially for challenging near-OOD cases, without requiring retraining or architectural changes.
Beyond label- and score-centric approaches, matrix approximation techniques have also been explored. SeTAR (Li et al., 2024b) applies selective low-rank approximation to CLIP’s weight matrices, aiming to retain the most informative components for OOD discrimination while filtering out spurious correlations. In a different vein, OT-DETECTOR (Liu et al., 2025a) formulates OOD detection as an optimal transport problem, quantifying the semantic and distributional discrepancy between test samples and ID labels, and further refining detection by semantic-aware content refinement.
Collectively, these developments illustrate a clear methodological evolution: from static post-hoc scoring and negative label augmentation, to adaptive proxy generation, label consistency modeling, prototype-based outlier exposure, and low-rank or optimal transport-based representations. This progression not only improves OOD detection accuracy but also enhances robustness in multi-label, semantically rich, and highly dynamic deployment scenarios. Furthermore, the interplay between proxy-based, consistency-based, and low-rank approximation methods suggests promising directions for unified frameworks that leverage the strengths of each paradigm.
@@
Recent works have sought to address these challenges by exploring the nuanced relationship between visual and semantic spaces. For instance, existing methods (Huang et al., 2024; Dai et al., 2023) observe that visually similar image categories may possess distinct properties in the semantic space, which can be exploited for more effective OOD detection. Dai et al. (2023) introduce a consistency-based calibration method that incorporates an object detector to mitigate hallucinations, thereby leveraging the world knowledge of LLMs for more accurate ID class descriptions. In contrast, Huang et al. (2024) propose ODPC, which generates peer classes based on both ID labels and OOD samples across textual and visual modalities. Notably, ODPC requires ID images for training an MLP, indicating that it is not a zero-shot approach. In terms of OOD scoring, Dai et al. (2023) adopt the MSP, while ODPC utilizes a KNN-based scoring strategy.
Building upon these foundations, subsequent studies have systematically investigated the behavior of LLMs in OOD detection. For example, Liu et al. (2024a) analyze the propensity of LLMs to detect near-OOD versus far-OOD samples, the impact of different fine-tuning strategies, and the suitability of various OOD scoring functions. Their findings reveal that LLMs are naturally adept at detecting far-OOD samples, that generative fine-tuning outperforms discriminative fine-tuning, and that cosine distance is an effective OOD score due to the anisotropic nature of LLM embedding spaces.
@@
Remark. As mentioned above, “zero-shot” here refers to no exposure to ID images and only access to ID labels. However,  Fort et al. (2021); Ding and Pang (2024) argue that in deployment, labels for OOD are either readily available. They propose a variant of the baseline approach that incorporates the names of OOD classes as candidate labels, resulting in improved performance. However, access to OOD labels is rare in real-life scenarios and seems less reasonable.
@@ 5.2. Few-shot OOD Detection Approaches
Studies. The direct way of adapting a large pre-trained model with several ID samples is fine-tuning.  Ming and Li (2024) and  Dong et al. (2023) conduct studies on the impact of fine-tuning on OOD detection within the context of VLMs, with a special focus on CLIP.  Ming and Li (2024) pay more attention to the role of parameter-efficient fine-tuning (PEFT) methods and the OOD score. The MCM score, introduced by  Ming et al. (2022b) as an innovative measure, alongside methods based on prompt learning for detecting OOD instances, is recognized for its effectiveness. Similarly,  Dong et al. (2023) conduct an extensive comparison between various fine-tuning approaches, such as PEFT and traditional methods, and discover that PEFT exhibits superior performance in detecting OOD instances, which echos with the conclusion of  Ming and Li (2024). Following this,  Kim et al. (2024) believe that Finetune Like You Pretrain (FLYP), a fine-tuning method, deserves attention due to its good performance on classification tasks. Specifically, FLYP is to mimic contrastive language-image pre-training as CLIP. After comparing the performance of FLYP and PEFT methods on zero-shot OOD detection,  Kim et al. (2024) find FLYP yields better OOD detection performance than PEFT. It should be noted that while  Fort et al. (2021) discuss “few-shot” OOD detection, their approach relies on a small number of outlier examples rather than ID samples, making it unsuitable for inclusion in this section.
@@
Few-shot OOD Detection. Few-shot OOD detection aims to leverage a limited amount of ID data to adapt LPMs, primarily through prompt learning or parameter-efficient fine-tuning. Early approaches such as CoOp (Zhou et al., 2022b) and ZegCLIP (Zhou et al., 2023) focus on optimizing the contextual words in prompts while keeping the backbone fixed, thus preserving the generalization ability of the pre-trained models. Building upon this, LoCoOp (Miyai et al., 2023a) and IDPL (Bai et al., 2024a) further enhance context vector learning from a near-ID perspective: LoCoOp maximizes entropy to separate ID-irrelevant local features (e.g., backgrounds) from textual embeddings of ID classes, while IDPL synthesizes ID-like outliers near the ID boundary and employs a diversity loss to encourage variance among sampled OOD candidates. In addition, MMFT (Kim and Hwang, 2025) projects image and text features into a shared hyperspherical space and introduces a cross-modal alignment loss function to promote the alignment of image and text representations in the hyperspherical space. During training, the loss function brings the image and text embeddings closer together while effectively distinguishing OOD texts from ID texts. However, these prompt-only methods may underutilize the rich information contained in image features. Recent advances such as GalLoP (Lafon et al., 2024) and Local-Prompt (Zeng et al., 2025) address this limitation by jointly optimizing global and local prompts. GalLoP leverages both global and region-level features via prompt dropout and region-level regularization to improve generalization and robustness, whereas Local-Prompt introduces orthogonal local prompt optimization to better exploit limited image data.
Beyond prompt learning, DSFG (Dong et al., 2023) highlights that conventional fine-tuning can erode critical OOD knowledge. To mitigate this, DSFG merges original and fine-tuned features before classifier training, thereby retaining broad OOD awareness and offering plug-and-play compatibility with other fine-tuning-based methods. Recently, SCT (Yu et al., 2024a) proposes a self-calibrated tuning framework that dynamically balances ID classification and OOD regularization by introducing modulation factors based on sample uncertainty, enabling more reliable adaptation with only a few ID samples. Complementary to this, GaCoOp (Tong et al., 2025) addresses the gradient conflict between ID classification and OOD regularization in prompt optimization. By decomposing and projecting gradients, GaCoOp aligns the optimization directions, thereby mitigating the adverse effects of background bias and improving few-shot OOD detection performance, especially when data is scarce.
Despite these advancements, overfitting remains a challenge due to the scarcity of labeled data. To tackle this, Dual-Adapter (Chen et al., 2024c) proposes a training-free few-shot method that constructs positive and negative adapters in a plug-and-play manner, requiring only a handful of ID samples. Meanwhile, to address the intrinsic modality gap in VLMs, SUPREME (Wang et al., 2025a) introduces multi-modal prototypes and alignment modules, while CLIP-OS (Sun et al., 2024) regularizes the ID/OOD boundary by synthesizing OOD data through mixing ID features, achieving strong performance without external OOD supervision. Additionally, NegPrompt (Li et al., 2024c) further refines the boundary by optimizing negative prompts, demonstrating the continued evolution and diversification of few-shot OOD detection strategies.
@@
Meta-learning based. Meta-learning aims to devise a learning approach that enables rapid adaptation to new challenges  (Vilalta and Drissi, 2002; Hospedales et al., 2021). OOD-MAML  (Jeong and Kim, 2020) adapt model-agnostic meta-learning (MAML) for few-shot OOD detection. It generates OOD samples and incorporates them along with ID data for the adapted N-way K-shot task, which is divided into N sub-tasks, each focusing on K-shot OOD detection. The decision on whether test data is OOD is based on the outcomes of these fast and simple N sub-tasks. In contrast, HyperMix  (Mehta et al., 2024) advocates for employing a hypernetwork-based method to enhance sample augmentation without the necessity for extra outliers. This is because classes not included in a specific meta-training task can act as OOD samples.
@@ 6.1. Evaluation metrics
AUPR (Area under the Precision-Recall curve). This metric is pertinent when the ID class is considered the positive class and is particularly valuable in the context of imbalanced class distributions. It assesses the balance between precision and recall, with a higher AUPR value indicating superior model performance. Therefore, a greater AUPR value is desirable.
@@
These metrics provide comprehensive insights into the OOD detection performance, considering different aspects such as the ability to distinguish between ID and OOD samples, handling imbalanced class sensitivity towards ID samples, and controlling the false positive rate at a specific true positive rate threshold.
@@ 6.2. Experimental Protocols
Recently,  Yang et al. (2023b); Bai et al. (2023); Yu et al. (2024b) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously. This is crucial to prevent the loss of model generalization. The samples mentioned earlier are termed as cs-ID data, an abbreviation for “covariate shift ID” data. Consequently, a new experimental protocol has been explored, called full-spectrum OOD detection. During the testing phase, the model is expected to identify near-OOD and far-OOD instances. Additionally, it should refuse to provide predictions for the OOD data and accurately predict ID and cs-ID data.
@@ 6.3. Application
To provide a comprehensive overview of recent advances in OOD detection, we first organize our discussion based on fundamental technical domains (Computer Vision, Natural Language Processing, and Beyond) as these domains have distinct methodological characteristics and theoretical foundations that shape how OOD detection is approached. This domain-based categorization allows us to systematically analyze the core technical innovations and challenges specific to each field. We then extend our analysis to an application-oriented perspective, as shown in Table 6.3, which bridges the gap between theoretical advances and practical implementations, illustrating how these domain-specific techniques are adapted and deployed in real-world scenarios.
@@
3D Object Detection Identifying out-of-distribution objects in 3D industrial environments, such as warehouses or robotics. (3D, ; Li et al., )
@@
Healthcare and Biology Medical Image Analysis Detecting unknown diseases or anomalies in medical images (e.g., X-ray, MRI). (Dis, ; Oh et al., 2024; Pokhrel et al., 2024; Smedsrud et al., 2021)
@@
Security and Surveillance Facial Authentication Detecting spoofed or unregistered faces for enhanced security. (Kahya et al., 2024)
@@
Audio Recognition Recognizing abnormal sounds such as alarms or intrusions. (Mazumder et al., 2021; Gong et al., 2022)
Scientific Research Astronomical Imaging Discovering unknown celestial objects or phenomena in astronomical data. (Zhou et al., 2024)
Solar Image Analysis Detecting novel solar activities or anomalies in solar observation images. (Giger and Csillaghy, 2024; Das et al., 2024)
Graph Neural Networks Detecting anomalies in complex scientific data such as molecular structures or social networks. (Bazhenov et al., 2022; Wu et al., 2023b; Jiang et al., 2025)
Mathematical Reasoning Identifying out-of-distribution problem types in automated reasoning systems. (Wang et al., 2024b)
@@
General-Purpose Applications Image Classification Detecting unknown categories in image classification systems. (Yang et al., 2023a, 2022; Sun et al., 2021)
@@ 6.3.1. Computer Vision
  • Image Classification. The majority of OOD detection research targets image classification, utilizing ID datasets such as MNIST (Xiao et al., 2017), CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009). Common OOD datasets include iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017), and Textures (Cimpoi et al., 2014). Full-spectrum OOD evaluation often employs the DIGITS, OBJECTS, and COVID benchmarks (Yang et al., 2023b). In addition to the aforementioned methods, most current research classifies OOD datasets based on the difficulty of distinguishing them from ID datasets. The most popular approach is to use ImageNet-1k as the ID dataset, with SSB-hard (Vaze et al., 2022) and NINCO (Bitterwolf et al., 2023) categorized as near-OOD (representing categories that are difficult to distinguish), and iNaturalist (Van Horn et al., 2018), Textures (Cimpoi et al., 2014), and OpenImage-O (Wang et al., 2022c) categorized as far-OOD (representing categories that are easier to distinguish
@@
  • Object Detection. OOD detection in object detection is relatively new (Du et al., 2022b), with evaluations on datasets like PASCAL-VOC (Everingham et al., 2015) and Berkeley DeepDrive-100K (Yu et al., 2018). Novel approaches include Situation Monitor, which uses diversity-driven budding ensemble architectures for zero-shot OOD detection in Transformer-based models (Syed et al., 2024), Proto-OOD, which enhances OOD object detection with prototype feature similarity (Chen et al., 2024d), and VisTa, which employs visual-contextual and text-augmented methods for zero-shot object-level OOD detection (Zhang et al., 2025b).
@@
  • Medical Image Analysis. OOD detection is vital in medical imaging, using datasets like CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021) depending on the image category (Tan et al., 2024). Recent studies focus on reliable OOD detection methods for digital pathology (Oh et al., 2024) and test-time augmentation for improving OOD detection in gastrointestinal vision (Pokhrel et al., 2024). In practice, IDx-DR (Dis, ), the first FDA-approved autonomous AI diagnostic system, implements OOD detection by rejecting images that fail to meet quality standards or contain anomalies, deferring these cases to human experts for assessment.
  • Human Action Recognition. OOD detection plays a crucial role in ensuring the robustness of Human Action Recognition (HAR) models, particularly when faced with previously unseen actions. By leveraging advanced techniques such as attention-based debiasing (Sim et al., 2023), generative feature synthesis (Mandal et al., 2019) and energy-based skeleton modeling (Xu et al., 2024b), recent research has significantly improved OOD detection performance. These methods are evaluated on benchmark datasets like UCF-101 (Soomro et al., 2012), HMDB-51 (Kuehne et al., 2011) and Kinetics-400 (Kay et al., 2017).
  • Solar Image Analysis. In space weather forecasting, OOD detection serves as a crucial tool for identifying solar anomalies. Recent unsupervised  (Giger and Csillaghy, 2024; Das et al., 2024)approaches leverage Solar Dynamics Observatory data to enhance detection accuracy.

### v4

@@ 5.1. Zero-shot OOD Detection Approaches
where $x’$ ’ denotes the test image, $Y_{in}$ is the set of ID labels, $T(t_{i})$ is the text embedding for prompt $t_{i}$ , $s_{i}(x’)$ ’ ) is the cosine similarity between the image and text features, and $\tau$ is the softmax temperature.
@@
Several works have enriched the textual context in CLIP-based OOD detection by augmenting candidate labels with curated corpora, as seen in NegLabel (Anonymous, 2024), LAPT (Zhang et al., 2024a), CSP (Chen et al., 2024b), and CLIPScope (Fu et al., 2024). Specifically, NegLabel selects negative labels from a corpus that are semantically dissimilar to ID classes, thereby sharpening the model’s ability to distinguish OOD samples. CLIPScope (Fu et al., 2024) goes a step further by integrating these negative labels into a Bayesian scoring framework to revise the original OOD score. LAPT, in contrast, utilizes text-to-image generation or image retrieval models to obtain visual exemplars for both ID and negative labels, followed by prompt tuning to optimize OOD discrimination. It is worth emphasizing that methods such as ZOC (Esmaeilpour et al., 2022), CLIPN (Wang et al., 2023c), NegLabel (Anonymous, 2024), and CLIPScope (Fu et al., 2024) each propose distinct OOD scoring mechanisms tailored to their respective architectures. In contrast, the MCM score can be seamlessly integrated into diverse CLIP-based frameworks.
Recent zero-shot OOD detection methods mainly differ in negative sampling, noise handling, and semantic grouping. AdaNeg (Zhang and Zhang, 2024) and AdaND (Cao et al., 2025) refine negative sampling and noise separation, while SimLabel (Zou et al., 2025) focuses on automatic semantic grouping. For outlier representation, OLE (Ding and Pang, 2024) expands outlier label diversity through clustering and interpolation, and prompt learning (Jung et al., 2024) introduces new scoring functions for post-hoc detection without retraining. Matrix-based methods further diversify the field, with SeTAR (Li et al., 2024b) applying low-rank approximations and OT-DETECTOR (Liu et al., 2025a) using optimal transport to capture semantic gaps.
@@
Recent works explore the relationship between visual and semantic spaces to improve OOD detection. Studies (Huang et al., 2024; Dai et al., 2023) leverage the distinct properties of visually similar categories in semantic space. Dai et al. (2023) develops a consistency-based method with object detection to enhance LLM-based class descriptions, while Huang et al. (2024) proposes ODPC to generate peer classes using both ID labels and OOD samples. These approaches differ in their scoring mechanisms, with Dai et al. (2023) using MSP and ODPC employing KNN-based scoring. Building upon these foundations, subsequent studies have systematically investigated the behavior of LLMs in OOD detection. For example, Liu et al. (2024a) analyze the propensity of LLMs to detect near-OOD versus far-OOD samples, the impact of different fine-tuning strategies, and the suitability of various OOD scoring functions.
@@
Remark. As mentioned above, “zero-shot” here refers to no exposure to ID images and only access to ID labels. While some works (Fort et al., 2021; Ding and Pang, 2024) improve performance by using known OOD class names as candidate labels, this approach relies on the often unrealistic assumption that OOD labels are readily available. This setting is also distinct from recent unsupervised fine-tuning approaches (Liang et al., 2024), which address a more realistic scenario where the unlabeled data for adaptation is possibly contaminated with OOD samples.
@@ 5.2. Few-shot OOD Detection Approaches
Studies. Fine-tuning large pre-trained models with limited ID samples is a common adaptation strategy. Recent studies (Ming and Li, 2024; Dong et al., 2023) systematically evaluate fine-tuning and parameter-efficient fine-tuning (PEFT) methods for OOD detection in VLMs, particularly CLIP. Both works highlight the superior OOD detection performance of PEFT over traditional fine-tuning. The MCM score (Ming et al., 2022b) and prompt-based approaches are also shown to be effective. FLYP (Kim et al., 2024), which mimics CLIP-style pretraining, further outperforms PEFT in zero-shot OOD detection. Notably, few-shot OOD detection using outlier examples (Fort et al., 2021) is not included here, as it does not align with the focus on ID samples.
@@
Few-shot OOD Detection. Few-shot OOD detection aims to leverage a limited amount of ID data to adapt LPMs, primarily through prompt learning or parameter-efficient fine-tuning. Early approaches such as CoOp (Zhou et al., 2022b) and ZegCLIP (Zhou et al., 2023) focus on optimizing the contextual words in prompts while keeping the backbone fixed, thus preserving the generalization ability of the pre-trained models. Building upon this, LoCoOp (Miyai et al., 2023a) and IDPL (Bai et al., 2024a) further enhance context vector learning from a near-ID perspective: LoCoOp maximizes entropy to separate ID-irrelevant local features (e.g., backgrounds) from textual embeddings of ID classes, while IDPL synthesizes ID-like outliers near the ID boundary and employs a diversity loss to encourage variance among sampled OOD candidates. In addition, MMFT (Kim and Hwang, 2025) projects image and text features into a shared hyperspherical space and introduces a cross-modal alignment loss function to promote the alignment of image and text representations in the hyperspherical space. However, these prompt-only methods may underutilize the rich information contained in image features. Recent advances such as GalLoP (Lafon et al., 2024) and Local-Prompt (Zeng et al., 2025) address this limitation by jointly optimizing global and local prompts.
Beyond prompt learning, recent methods address the limitations of conventional fine-tuning, such as the loss of OOD awareness, by merging original and adapted features (Dong et al., 2023) or introducing self-calibrated tuning frameworks that balance ID and OOD objectives (Yu et al., 2024a). Other approaches improve prompt optimization through gradient-based techniques (Tong et al., 2025). To mitigate overfitting under limited supervision, strategies such as plug-and-play adapters (Chen et al., 2024c), multi-modal alignment (Wang et al., 2025a), and boundary regularization with synthesized OOD features (Sun et al., 2024) have been proposed. Further refinements include optimizing negative prompts to better distinguish ID/OOD boundaries (Li et al., 2024c). Collectively, these methods enhance few-shot OOD detection by improving adaptation, alignment, and boundary modeling.
@@
Meta-learning based. Meta-learning aims to devise a learning approach that enables rapid adaptation to new challenges  (Hospedales et al., 2021). OOD-MAML  (Jeong and Kim, 2020) adapt model-agnostic meta-learning (MAML) for few-shot OOD detection. It generates OOD samples and incorporates them along with ID data for the adapted N-way K-shot task, which is divided into N sub-tasks, each focusing on K-shot OOD detection. The decision on whether test data is OOD is based on the outcomes of these fast and simple N sub-tasks. In contrast, HyperMix  (Mehta et al., 2024) advocates for employing a hypernetwork-based method to enhance sample augmentation without the necessity for extra outliers. This is because classes not included in a specific meta-training task can act as OOD samples.
@@ 6.1. Evaluation metrics
AUPR (Area under the Precision-Recall curve). This metric is pertinent when the ID class is considered the positive class and is particularly valuable in the context of imbalanced class distributions. It assesses the balance between precision and recall, with a higher AUPR value indicating superior model performance.
@@ 6.2. Experimental Protocols
Recently,  Yang et al. (2023b); Bai et al. (2023) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously. This is crucial to prevent the loss of model generalization. The samples mentioned earlier are termed as cs-ID data, an abbreviation for “covariate shift ID” data. Consequently, a new experimental protocol has been explored, called full-spectrum OOD detection. During the testing phase, the model is expected to identify near-OOD and far-OOD instances. Additionally, it should refuse to provide predictions for the OOD data and accurately predict ID and cs-ID data.
@@ 6.3. Application
To comprehensively review recent advances in OOD detection, we first categorize our discussion by key technical domains (Computer Vision, Natural Language Processing, and Beyond), as each has distinct methods and theoretical bases shaping OOD approaches. This structure enables a systematic analysis of core innovations and challenges in each field. We then adopt an application-oriented perspective (see Table 6.3), highlighting how these techniques are adapted and implemented in real-world scenarios.
@@
3D Object Detection Identifying OOD objects in 3D industrial environments, such as warehouses or robotics. (3D, ; Li et al., )
@@
Healthcare and Biology Medical Image Analysis Detecting unknown diseases or anomalies in medical images (e.g., X-ray, MRI). (Dis, ; Pokhrel et al., 2024; Smedsrud et al., 2021)
@@
Securit and Surveillance Facial Authentication Detecting spoofed or unregistered faces for enhanced security. (Kahya et al., 2024)
@@
Audio Recognition Recognizing abnormal sounds such as alarms or intrusions. (Mazumder et al., 2021)
Scientifi Research Astronomical Imaging Discovering unknown celestial objects or phenomena in astronomical data. (Zhou et al., 2024)
Solar Image Analysis Detecting novel solar activities or anomalies in solar observation images. (Das et al., 2024)
Graph Neural Networks Detecting anomalies in complex scientific data such as molecular structures or social networks. (Bazhenov et al., 2022; Wu et al., 2023b)
Mathematical Reasoning Identifying OOD problem types in automated reasoning systems. (Wang et al., 2024b)
@@
General-Purpos Applications Image Classification Detecting unknown categories in image classification systems. (Yang et al., 2023a, 2022; Sun et al., 2021)
@@ 6.3.1. Computer Vision
  • Image Classification. Image classification represents the primary testbed for OOD detection research. Common ID datasets include MNIST (Xiao et al., 2017), CIFAR (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009), while OOD evaluation typically uses datasets like iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), and Textures (Cimpoi et al., 2014). Recent research categorizes OOD datasets based on their detection difficulty: near-OOD (e.g., SSB-hard (Vaze et al., 2022), NINCO (Bitterwolf et al., 2023)) for harder-to-distinguish categories, and far-OOD (e.g., iNaturalist, OpenImage-O (Wang et al., 2022c)) for more distinct categories.
@@
    Object Detection. OOD detection for object detection tasks is a relatively recent research area (Du et al., 2022b), with evaluations commonly conducted on datasets such as PASCAL-VOC (Everingham et al., 2015) and BDD-100K (Yu et al., 2018). Recent methods explore diverse strategies, including ensemble architectures (Syed et al., 2024), prototype-based similarity (Chen et al., 2024d), and visual-contextual augmentation (Zhang et al., 2025b).
@@
  • Medical Image Analysis. OOD detection is vital in medical imaging, using datasets like CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021) depending on the image category (Tan et al., 2024). Recent studies focus on reliable OOD detection methods for digital pathologyfor improving OOD detection in gastrointestinal vision (Pokhrel et al., 2024). In practice, IDx-DR (Dis, ), the first FDA-approved autonomous AI diagnostic system, implements OOD detection by rejecting images that fail to meet quality standards or contain anomalies, deferring these cases to human experts for assessment.
  • Human Action Recognition. OOD detection plays a crucial role in ensuring the robustness of Human Action Recognition (HAR) models, particularly when faced with previously unseen actions. By leveraging advanced techniques such as attention-based debiasing (Sim et al., 2023), generative feature synthesis (Mandal et al., 2019) and energy-based skeleton modeling (Xu et al., 2024b), recent research has significantly improved OOD detection performance.
  • Solar Image Analysis. In space weather forecasting, OOD detection serves as a crucial tool for identifying solar anomalies. Recent unsupervised  (Das et al., 2024)approaches leverage Solar Dynamics Observatory data to enhance detection accuracy.